Consider the demand for a good. At price Rs 4, the demand for the good is 25 units.

Suppose price of the good increases to Rs 5, and as a result, the demand for the good

falls to 20 units. Calculate the price elasticity?


1
Expert's answer
2021-09-22T03:47:02-0400

Price Elasticity=%change in quantity/%change in price

Percentage Change in Quantity=[(20 units*100)/25 units]-100=-20%

Percentage Change in Price= [(Rs4+ Rs5)*100/4]-100=125%

Therefore price elasticity for demand =-20/125

=0.16